Solve the following simultaneous equations by using Cramer’s rule


3x+2y=13
2x-y=4

Answer by mananth(16946)

3x+2y=13
2x-y=4
d=| 3 2 |
...| 2 -1|

= 3*-1 -2*2
=-7
Dx= | 13 2|
... | 4 -1|
=-13-8
=-21

Dy= | 3 13|
... | 2 4|
= 12-26
=-14
x=Dx/D = -21/-7
x=3
y=Dy/D = -14/-7=2
x=3, y=2